OpenStudy (anonymous):

what is the value of cot7 1/2 degree? without using calculator !

OpenStudy (cwrw238):

cot x = 1 / tan x use your calculator to find tan 7.5 degrees and plug this into the above

OpenStudy (anonymous):

well i have some options

OpenStudy (anonymous):

a) sq root 6+ sq r 2 + sq r 3 + sq r 4 b) sq r 6 + sq r 2 - sq r 3 - sq r 4 c) sq r 6 -sqr 2- sq r 3- sqr 4 d sq r 6 + sq r 2 - sq r 3- sqr 4

OpenStudy (anonymous):

is that cot(7.5)? @kritima @cwrw238 the question says you can't use a calculator =/

OpenStudy (anonymous):

yeh it is cot 7.5

OpenStudy (anonymous):

I don't see how you would do this without a calculator though lol.

OpenStudy (anonymous):

i know that tan 7.5 degree is ( sq r 3- sq r 2)/ ( sq r 2 - 1 )

OpenStudy (anonymous):

so to find cot(7.5) then that's just 1/tan(7.5). Since you know what tan(7.5) is, to get 1/tan(7.5), just flip the fraction and you'll have the answer. @kritima

OpenStudy (anonymous):

i did that but i didn't get the answer !

OpenStudy (anonymous):

my answer isn't in the option

OpenStudy (anonymous):

Oh that's because you didn't rationalise the denominator. So here is what you need to do. We know the following information:\[\bf \tan(7.5)=\frac{\sqrt{3}-\sqrt{2}}{\sqrt2-1}\]Now to find cot(7.5), we just plug that in:\[\bf \cot(7.5)=\frac{1}{\tan(7.5)}=\frac{ \sqrt2-1 }{ \sqrt3-\sqrt2 }\]Now we must rationalise the denominator, i.e. we must eliminate the roots in the denominator. We do this by multiplying both the numerator and denominator of the fraction by the conjugate. The conjugate is \(\bf \sqrt{3}+\sqrt{2}\), as you see we just changed the minus to a plus.\[\bf \frac{ \sqrt2-1 }{ \sqrt3-\sqrt2 }=\frac{ (\sqrt2-1)(\sqrt3+\sqrt2) }{ (\sqrt3-\sqrt2)(\sqrt3+\sqrt2) }=\frac{\sqrt{6}+\sqrt4-\sqrt{3}-\sqrt2}{1}\]Can you tell me what the answer will be now? @kritima

OpenStudy (anonymous):

cot(7.5) is 1/tan(7.5) then 7.5 is the half-angle of 15...which in turn is the half angle of 30. so it's a bunch of half-angle formulas. i assume that's what you did?

OpenStudy (anonymous):

i also got the same @genius12

OpenStudy (anonymous):

@kritima I think the choices are typed up wrong?

OpenStudy (anonymous):

but if you use half angle formula over here then you'll get completly different answer !

OpenStudy (anonymous):

@kritima forget that, I'm saying that perhaps the choices are typed up wrong because I'm pretty confident that's the right answer.

OpenStudy (anonymous):

i aslo thought so @genius12

OpenStudy (anonymous):

but when you use half angle formula you'll get one which is in option

OpenStudy (anonymous):

so i m bit confused which is the right answer

OpenStudy (anonymous):

@kritima I see what the problem is now. The given value of tan(7.5) is incorrect. Are you that is what is given?

OpenStudy (anonymous):

@genius12 sadly, the answer you're confident in does not match a calculator's answer. you may not use a calculator in answering, but you may use it in checking after all

OpenStudy (anonymous):

@abb50 I was assuming that the given value of tan(7.5) is correct, which is the one @kritima gave me. Turns out the value he gave me was wrong which is why the answer is wrong.

OpenStudy (anonymous):

although, one of the choices matches the calculator's answer

OpenStudy (anonymous):

ok i made mistake the value of tan 7.5 i gave you was wrong @ genius12 . i m sorry !

OpenStudy (anonymous):

tan(7.5) is sqrt 2 + 1 not minus..

OpenStudy (anonymous):

now i got it all right . the answer is option a

OpenStudy (anonymous):

yeh you are right @abb50

OpenStudy (anonymous):

@kritima why did you give me the wrong value, u said sqrt(2) - 1 lol...

OpenStudy (anonymous):

actully i made a mistake. and i was also assuming the same sorry ! @genius12

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

and thank you all of you !